# Observing dynamics of distinct structural transitions in trapped-ion clusters

### v2

@@ III.3 Hysteresis and a triple point
The vastly unequal regions of stability of the two configurations (pentagon vs pyramid) lead to a hysteresis when $\alpha$ is tuned back and forth across the transition point, particularly beyond the spinodal point at $\alpha=1.29$. We apply triangular voltage ramps for $U_{dc}$ (see Fig. 1) to observe this effect in the experiment. As shown in Fig. 4(c), the pentagon sharply transitions to a square pyramid close to $\alpha_{t}$ during the axial decompression phase. However, when the cluster is axially compressed by increasing $\alpha$, the system gets trapped in the local minimum, and the pyramidal arrangement remains intact beyond the transition point. The transition back to the pentagon configuration occurs significantly beyond $\alpha_{t}$. This phenomenon is analogous to the well-known example of metastable super-cooled water, which freezes via nucleation or an external disturbance (in this case, a photon kick or background collision). A few ramps during the compression phase are shown to demonstrate this nucleation effect in the transition back to the pentagon, which occurs over a large range of $\alpha$. Two real-time experimental videos demonstrating fast and slow nucleation are provided as supplementary material.
IV Discussion and Conclusion
We directly observe dynamical signatures of distinct classes of structural transitions in a unit-cell-sized Coulomb cluster of laser-cooled ions in an rf trap. The excellent agreement between the experimentally observed dynamics, analytical calculations based on symmetry of the configurations, and numerically computed collective-mode spectra provides a comprehensive picture of different transition mechanisms. Continuous transitions exhibit mode softening, where the frequency of the symmetry-breaking mode vanishes at the transition point. This effect is visualized by measuring the response of a $4$-ion cluster to a controlled trap modulation. Second, a triple point-like feature is observed in a $5$-ion cluster, where a pentagonal ground state undergoes a symmetry-breaking transition at the same point where the global minimum jumps from the pentagon to a pyramid with a different symmetry. We observe hysteresis by cyclically tuning the confinement and tracking the structural memory of the $5$-ion cluster. Finally, at a discontinuous transition in a $6$-ion cluster, the configuration switches stochastically between two distinct symmetry groups.
@@ IV Discussion and Conclusion
Previously, such switching behavior was observed in ion chains in a regime where background gas collisions are the dominant mechanism, enabling the characterization of ultra-high vacuum pressures [39, 40]. The ability to realize and monitor such transitions in a 3D unit cell with tunable energy barrier will enable the study of thermal activation processes due to photon kicks from the Doppler-cooling laser. Furthermore, cooling the crystal to its motional ground state would enable studies of the structural transitions in the quantum regime [41, 42, 33]. One can also employ such finite, self-organized clusters to investigate thermalization and heat transport [43, 44], pattern formation [45], crystals with impurities and defects [46, 47, 48, 49, 50] and Floquet engineering [51] in a controlled, strongly-correlated environment, broadening our understanding of collective phenomena in the presence of long-range interactions.
Acknowledgements.
We thank Arun Roy for comments on manuscript. The authors acknowledge support from Department of Science and Technology and Ministry of Electronics and Information Technology (MeitY), Government of India, under the “Centre for Excellence in Quantum Technologies” grant with Ref. No. 4(7)/2020-ITEA. A. Paramekanti acknowledges funding from the Natural Sciences and Engineering Research Council (NSERC) of Canada.
Supplementary discussion
I End-cap Trap
@@
The time-varying potential generated by the trap is of the form:
$\Phi(\bm{r},t)=\eta(U_{dc}+V_{rf}\cos(\Omega_{rf}t))\times\frac{x^{2}+y^{2}-2z^{2}}{4z_{0}^{2}},$ (S1)
@@
where $z_{0}$ is half the distance between rf-electrodes and $\eta$ is the quadrupole efficiency of the trap geometry [1, 2]. The dc ($U_{dc}$) and rf ($V_{rf}$) voltages are set to operate the trap within the first Mathieu stability region [3], represented in the rescaled Mathieu a ($\propto U_{dc}$) and q ($\propto V_{rf}$) parameter space. Under the adiabatic approximation - [4] $|a|,q^{2}\ll 1$, the angular frequencies of the time-averaged harmonic potential are related to the Mathieu parameters as:
$\omega_{x,z}$ $=$ $\frac{\Omega_{rf}}{2}\sqrt{\frac{q_{x,z}^{2}}{2}+a_{x,z}},$ (S2)
@@
where $a_{z}=-2a_{x}=\frac{4\eta QU_{dc}}{mz_{0}^{2}\Omega_{rf}^{2}}$, $q_{z}=-2q_{x}=\frac{-2\eta QV_{rf}}{mz_{0}^{2}\Omega_{rf}^{2}}$, Q is the charge of the trapped particles ($Q=+e$ for singly ionized Calcium), and $\Omega_{rf}=18.26$ MHz is the drive frequency of the trap. By fixing $V_{rf}\sim 235$ V and varying $U_{dc}\sim 1$ V to $-12$ V, $\alpha=\omega_{z}/\omega_{x}$ can be continuously tuned from $\sim 2.2$ to $0.25$. The fluorescence images of the clusters are imaged onto an EMCCD camera using a custom-made objective with $\times$22 magnification.
II Exciting the soft mode of the 4-ion tetrahedron
@@
The motion of trapped ions in our experiment is excited via two different methods. The first method is forced oscillation of the trap center, where one of the conical DC electrodes is given a small sinusoidal voltage, which results in the center-of-mass (c.o.m.) of the cluster oscillating about the trap center. When this frequency matches the c.o.m. mode’s frequency, the increase in the oscillation amplitude of the ions is seen on EMCCD. Any slight misalignment of the conical electrode that is inadvertent leads to an oscillatory gradient field along both the radial and axial directions, enabling the excitation of both c.o.m. modes on resonance with this scheme. We directly measure $\alpha$ at different $U_{dc}$ values by exciting a single ion’s radial and axial motion.
The second method is the amplitude modulation of rf-voltage ($V_{rf}$) at a frequency $\Omega_{m}$. This leads to an oscillatory Mathieu $q$ parameter, effectively modulating the elastic constants of the time-averaged harmonic potential, $\omega_{x}^{2}$ and $\omega_{z}^{2}$. This leads to a parametric excitation of the c.o.m. motion for $\Omega_{m}=2\times\omega_{x,z}$ [5]. In the case of a multi-ion cluster, the modulation of trap curvatures will also drive the ions about the mean CME coordinate $\bm{v}_{0}$ along the unit vector $\bm{v}_{m}$.
Figure SF1: Toy-problem of 2 ions trapped in 1D harmonic potential. The two normal modes are c.o.m. oscillation (blue arrows) and breathing mode (green arrows). Modulation of the trap curvature leads to the driven motion along the red arrows. The Fourier spectrum of the motion shows on-resonant excitation of the breathing mode via curvature modulation.
To obtain this unit vector at a set of ($V_{rf}$, $U_{dc}$) values, the CME coordinates $\bm{v}_{0}^{\pm\delta}$ corresponding to small perturbation in rf-voltage $V_{rf}(1\pm\delta)$ are computed. For small values of $\delta$ $(\sim 10^{-3})$, we obtain
$\bm{v}_{m}=\frac{\bm{v}_{0}^{+\delta}-\bm{v}_{0}^{-\delta}}{||\bm{v}_{0}^{+\delta}-\bm{v}_{0}^{-\delta}||},$ (S3)
the unit vector along the driven motion about $\bm{v}_{0}$. The collective modes that can couple to the drive can be obtained from the projections $|\bm{v}_{m}\cdot\bm{u}_{k}|$, where $\bm{u}_{k}$ are the normal-mode eigen vectors.
@@
Table 1: Projection of the driven motion onto the 12 normal-mode eigenvectors of the $4$-ion cluster.
@@
To illustrate the resonant excitation of a collective mode via trap curvature modulation, consider a toy problem of two like charges with mass m, co-trapped in a 1D harmonic potential of frequency $\omega_{t}=\sqrt{2k/m}=2\pi\times 1$ MHz. The two normal modes are shown in Fig. SF1: c.o.m. mode with frequency $\omega_{c.o.m.}=\omega_{t}$ and breathing mode with frequency $\omega_{b}=\sqrt{3}\times\omega_{t}$. The motion of the charges is first damped to equilibrium positions, and the spring constant $k$ is then modulated at $\Omega_{m}$, i.e., $k\rightarrow k(1+\epsilon\cos(\Omega_{m}t))$. In Fig. SF1, the power spectrum of the ions’ motion after the damping period is compared with the power spectra after the modulation at different frequencies. For a modulation strength $\epsilon=10^{-3}$, the c.o.m. mode shows parametric excitation at $\Omega_{m}=2\times\omega_{t}$, while the breathing mode shows resonant excitation for $\Omega_{m}=\sqrt{3}\times\omega_{t}$. In this case, it is trivial to see that the unit vector along the driven motion is exactly the breathing mode (Fig. SF1).
@@
Let us look at the collective modes of the $4$-ion cluster near the square to tetrahedron transition. The projections of all the eigenvectors along the drive are tabulated in Table 1. The symmetry-restoring soft mode of the tetrahedron (shown in Fig. SF2) has a projection $\sim 1$, enabling the resonant excitation of this mode experimentally. However, the same technique excites a high-frequency in-plane mode of the square shown in Fig. SF2, and cannot be used to observe the softening of the square’s symmetry-breaking mode.
Figure SF2: (Top) Soft-mode eigenvectors of square and tetrahedron. (Bottom) The driven motion due to trap voltage modulation. The arrows with the same color represent the direction of the in-phase drive for each ion.
III Mode softening of the 5-ion pentagon
III.1 Out-of-plane normal modes
@@
where $l$ is an integer characterizing the angular dependence. These modes are sketched in Fig. SF3: c.o.m. Z-Motion for Cosine l=0 mode, rigid tilt/rotation about Y-axis for Cosine l=1 mode, rigid tilt/rotation about X-axis for Sine l=1 mode, and two buckling modes for Cosine and Sine l=2.
III.2 Buckling critical point

### v3

@@ II.3 Hysteresis and a triple point
By applying triangular voltage ramps for $U_{dc}$ (see Fig. 1), we drive the $5$-ion cluster back and forth in a series of axial compression and decompression cycles. In each ramp, as the axial confinement is weakened, the pentagon sharply transitions to a square pyramid beyond $\alpha_{t}$ with the apical ion in the pyramid randomly appearing at $z>0$ or $z<0$. This is signaled by the sudden increase in $|\psi_{30}|$ with decreasing $\alpha$ (rightward sweep) as shown in Fig. 3(b). However, when the cluster is axially compressed by increasing $\alpha$ (leftward sweep), the pyramidal structure remains intact well beyond $\alpha_{t}$, as the system remains trapped in a local minimum. This phenomenon is analogous to the well-known example of metastable supercooled water. A photon kick or background collision triggers a nucleation-like event into the pentagonal ground state. These events occur over a broad range of $\alpha$, where the initial finite $|\psi_{30}|$ randomly jumps to zero. The resulting hysteresis loop is strongly asymmetric as seen from Fig. 3(b). Real-time experimental videos demonstrating fast and slow nucleation events are provided as supplementary material.
II.4 Stochastic switching
In the $6$-ion cluster, a reconstructive transition is observed between two 3D structures with different symmetries: A pentagonal pyramid with one apex ion on the $z$-axis switches to an octahedron (square bipyramid) with the two apex ions on the $z$-axis, as shown in Fig. 4(a). The lowest-order multipole that distinguishes these configurations is again the real, parity-odd octupole moment $\psi_{30}$, which vanishes for the octahedron and takes on positive (negative) values for the pyramid with the apical ion at $z>0$ ($z<0$).
The energy landscape obtained via the NEB method is shown in Fig. 4(a). The octahedron emerges as a stable node via subcritical pitchfork bifurcation at $\alpha\!\approx\!1.46$. The global minimum shifts between the pentagonal pyramid and the octahedron states at $\alpha\approx 1.238$ (Supplementary discussion I), and is also marked by a discontinuous jump in the normal-mode spectra (see SF6). The two pyramids vanish from the potential landscape at $\alpha\!\approx\!1.19$ via saddle-node bifurcations. The extended region of tri-stability between the octahedron and the two pentagon-pyramids leads to stochastic switching of the cluster configuration.
We experimentally track the dynamics of $\psi_{30}$ close to the reconstructive transition point, where all the minima have equal energy. Fig. 4(b) shows the time trace of $\psi_{30}$ obtained experimentally at $\alpha\!\approx\!1.24$, which exhibits the expected three-state telegraph-like switching dynamics. We provide the video corresponding to the time trace as supplementary material.
III Discussion
In summary, we investigate three distinct dynamical signatures of structural transitions in unit-cell-like 3D trapped-ion clusters. While our system lacks the ability to address individual ions, the diverse octupolar charge orders and rich energy landscapes that we present open avenues for further fundamental studies. By cooling the system to the motional ground state [48], structural transitions can be studied in the quantum regime [49, 50, 51, 52]. By introducing spin-spin interactions [16, 53, 18] between ions, one can explore frustrated dynamics arising from competing charge and spin orders [54, 55, 19]—a frontier in many-body quantum physics.
@@ III Discussion
More readily, the stochastic switching between different cluster configurations can be applied to study reaction kinetics. Previously, such switching behavior was observed in zig-zag ion chains in a regime where background gas collisions are the dominant mechanism [56, 57]. The ability to tune the barrier heights and noise sources (e.g., laser photon recoil) in 2D/3D energy landscapes [58] provides an ideal testbed for studies of thermal activation processes. One can also employ these controllable 2D and 3D clusters to investigate stochastic resonance [59], pattern formation [60], heat transport and thermalization [61, 62, 63], dynamics in the presence of impurities [64] and topological defects [65, 66, 67, 68], as well as Floquet engineering [69], broadening our understanding of collective nonlinear phenomena in the presence of long-range interactions.
IV Materials and Methods
IV.1 Experimental system
@@
The time-varying potential generated by the end-cap type trap shown in Fig. 1 is of the form:
Figure 5: (Top) Soft-mode eigenvectors of square and tetrahedron. (Bottom) The arrows with the same color represent in-phase motion of each ion due to trap voltage modulation.
$\Phi(\bm{r},t)=\eta(U_{dc}+V_{rf}\cos(\Omega_{rf}t))\times\frac{x^{2}+y^{2}-2z^{2}}{4z_{0}^{2}},$ (2)
@@
where $z_{0}$ is half the distance between rf-electrodes in the axial direction and $\eta$ is the quadrupole efficiency of the trap geometry [70, 71]. The dc ($U_{dc}$) and rf ($V_{rf}$) voltages are set to operate the trap within the first Mathieu stability region [72], represented in the rescaled Mathieu a ($\propto U_{dc}$) and q ($\propto V_{rf}$) parameter space. Under the adiabatic approximation [73], $|a|,q^{2}\ll 1$, the angular frequencies of the time-averaged harmonic potential are related to the Mathieu parameters as:
$\omega_{x,z}$ $=$ $\frac{\Omega_{rf}}{2}\sqrt{\frac{q_{x,z}^{2}}{2}+a_{x,z}},$ (3)
@@
where $a_{z}=-2a_{x}=\frac{4\eta QU_{dc}}{mz_{0}^{2}\Omega_{rf}^{2}}$, $q_{z}=-2q_{x}=\frac{-2\eta QV_{rf}}{mz_{0}^{2}\Omega_{rf}^{2}}$, Q is the charge of the trapped particles ($Q=+e$ for singly ionized Calcium), and $\Omega_{rf}=18.26$ MHz is the drive frequency of the trap. By fixing $V_{rf}\sim 235$ V and varying $U_{dc}\sim 1$ V to 12-12 V, $\alpha=\omega_{z}/\omega_{x}$ can be continuously tuned from $\sim 2.2$ to $0.25$. The fluorescence images of the clusters are captured onto an EMCCD camera with a custom-made objective of $\times$22 magnification.
IV.2 Trapped ion motional excitation
@@
We excite the motion of trapped ions in the experiment via two different methods. The first method is forced oscillation of the trap center, where one of the conical DC electrodes is given a small sinusoidal voltage, which results in the center-of-mass (c.o.m.) of the cluster oscillating about the trap center. When this frequency matches the c.o.m. mode’s frequency, the increase in the oscillation amplitude of the ions is seen on EMCCD. Any slight misalignment of the conical electrode that is inadvertent leads to an oscillatory gradient field along both the radial and axial directions, enabling the excitation of both c.o.m. modes on resonance with this scheme. We directly measure $\alpha$ at different $U_{dc}$ values by exciting a single ion’s radial and axial motion. The second method is the amplitude modulation of rf-voltage ($V_{rf}$) at a frequency $\Omega_{m}$. This leads to an oscillatory Mathieu $q$ parameter, effectively modulating the elastic constants of the time-averaged harmonic potential, $\omega_{x}^{2}$ and $\omega_{z}^{2}$. This provides an alternative scheme for parametric excitation of the c.o.m. motion for $\Omega_{m}=2\times\omega_{x,z}$ [74].
@@
Table 1: Projections of the driven motion due to voltage modulation onto the 12 normal-mode eigenvectors of the $4$-ion cluster.
@@
However, in the case of a multi-ion cluster, the modulation of trap curvatures will also drive the ions about the equilibrium configuration $\bm{v}_{0}$ along the unit vector $\bm{v}_{m}$. We obtain this unit vector at a set of ($V_{rf}$, $U_{dc}$) values, by numerically computing the CME coordinates $\bm{v}_{0}^{\pm\delta}$ corresponding to small perturbation in rf-voltage $V_{rf}(1\pm\delta)$. For small values of $\delta$ $(\sim 10^{-3})$,
[…]
Collective normal modes (eigen vectors $\bm{u}_{k}$) that can couple to this drive have a finite projection $|\bm{v}_{m}\cdot\bm{u}_{k}|$.
@@
Let us look at the effect of trap voltage modulation on the collective modes of $4$-ion cluster near the square to tetrahedron transition. The projections between the drive and the normal modes are tabulated in Table 1. The symmetry-restoring soft mode of the tetrahedron (shown in Fig. 5) has a projection $\sim 1$, enabling the resonant excitation of this mode experimentally. The same technique, however, excites a high-frequency in-plane mode of the square shown in Fig. 5, and cannot be used to observe the softening of the square’s symmetry-breaking mode. To demonstrate the resonant normal mode excitation via curvature modulation, we present a simple toy problem of two ions in a one-dimensional harmonic potential in the supplementary discussion III. In this simple case, the modulation of the spring constant leads to an on-resonant excitation of the breathing mode.
Figure 6: Heat maps showing Abel transforms of the $6$-ion pyramid (top) and the octahedron (bottom) obtained from fluorescence images to calculate the octupole moment of the charge distribution.
IV.3 Numerical computations of the energy landscape
The equilibrium positions of the ions can be efficiently obtained by minimizing the total potential energy (eq. 1) about a guess configuration, using the built-in FindMinimum function of Mathematica. For each $\alpha$ value, the initial configuration of the cluster is randomized and then steered to a local minimum. This is repeated several times to identify the global minimum (CME). The collective-mode eigenvectors ($\bm{u}_{k}$) and their corresponding frequencies ($\omega_{k}$) at the theoretical limit $T=0$ are then obtained from the Hessian matrix
where $v_{i}$ are the components of the vector $\bm{v}=(x_{1},y_{1},z_{1},...,x_{n},y_{n},z_{n})$ and $\bm{v}_{0}$ is the coordinate of the equilibrium configuration (local or global minimum).
To obtain the minimum energy path (MEP) and the saddle point between two equilibrium configurations, we employ a chain of states method. A guess path between the stable equilibria is constructed using linear interpolation. The entire chain is taken to the MEP with the NEB method [46]. Around the configuration with maximum energy in the MEP, we find an unstable equilibrium in the potential energy landscape, with one negative eigenvalue of its corresponding Hessian. The contour plots of energy landscapes are obtained by stitching the MEPs between consecutive minima, sorted by a suitable order parameter.
IV.4 Computing moments from images
The presence of cylindrical symmetry about z-axis in our system permits us to apply Abel transforms [75, 76] on the fluorescence images to reconstruct a time-averaged 3D charge distribution. Fig. 6 shows the Abel transforms obtained for the $6$-ion 3D orientations—pentagonal pyramid and octahedron. We use this 3D charge distribution to compute the octupole moments $\psi_{30}$ for the $5$-ion hysteretic dynamics, and the $6$-ion stochastic switching.
Acknowledgements
We thank Arun Roy for comments on the manuscript. The authors acknowledge support from the Department of Science and Technology and the Ministry of Electronics and Information Technology (MeitY), Government of India, under the “Centre for Excellence in Quantum Technologies” grant with Ref. No. 4(7)/2020-ITEA. A. Paramekanti acknowledges funding from the Natural Sciences and Engineering Research Council (NSERC) of Canada.
Supplementary discussion
I Analytical calculations of transition points
For the ease of calculating the transition points, the energy and length can be scaled into dimensionless units as $V\rightarrow V/E_{\perp}$ and $\bm{r}\rightarrow\bm{r}/l_{\perp}$, where
$l_{\perp}=\left(\frac{k_{e}e^{2}}{m\omega_{x}^{2}}\right)^{1/3}\quad\text{and}\quad E_{\perp}=\left(m\omega_{x}^{2}k_{e}^{2}e^{4}\right)^{1/3}.$ (S1)
Figure SF1: The CMEs of 4-, 5-, and 6-ion clusters at the transition points of interest.
I.1 4-ion square to tetrahedron
In these scaled units, the potential energy of the tetrahedral configuration in terms of the coordinates $\rho$, $z$ (depicted in Fig. SF1) is
$V_{4}(\rho,z)=2\rho^{2}+2\alpha^{2}z^{2}+\frac{1}{\rho}+\frac{4}{\sqrt{2\rho^{2}+4z^{2}}},$ (S2)
Simultaneously solving $\partial V_{4}/\partial\rho=0$ and $\partial V_{4}/\partial z=0$ for the equilibrium positions $(\rho^{0},z^{0})$, we obtain
$z^{0}=\begin{cases}0&\alpha\geq\alpha_{c}\\ \frac{1}{2}\,\left[\frac{1}{4}\left(\frac{2}{\alpha^{2}}\right)^{2/3}-\frac{1}{2}\left(\frac{1}{4(2-\alpha^{2})}\right)^{2/3}\right]^{1/2}&\alpha<\alpha_{c}\end{cases}$ (S3)
where $\alpha_{c}=2\sqrt{\frac{4-\sqrt{2}}{7}}\approx 1.216$ is the critical point where $z^{0}$ becomes non-zero, breaking inversion symmetry.
I.2 5-ion pentagon to square-pyramid
The potential energies in scaled units of the pentagonal and the square-pyramidal configurations are, respectively,
$V^{pentagon}_{5}(\rho_{1})=$ $\;\frac{5}{2}\rho_{1}^{2}+\frac{\sqrt{25+10\sqrt{5}}}{\rho_{1}},$ (S4)
$V^{pyramid}_{5}(\rho_{2},z)=$  $\;2\rho_{2}^{2}+10\alpha^{2}z^{2}+\frac{1+2\sqrt{2}}{\rho_{2}}+\frac{4}{\sqrt{\rho_{2}^{2}+25z^{2}}},$ (S5)
where the coordinates $\rho_{1}$, $\rho_{2}$, and $z$ are depicted in Fig. SF1. Solving for the equilibrium configurations $\rho_{1}^{0}$ and $(\rho_{2}^{0},z^{0})$, respectively, the minimum energy of the pentagon is obtained as $V^{pentagon}_{5}(\rho_{1}^{0})\approx 9.28$, while the minimum energy of the square-pyramid decreases with $\alpha$. Both curves intersect at $\alpha_{t}\approx 1.328$ as shown in Fig. SF1.
I.3 6-ion pentagonal-pyramid to octahedron
The potential energies in scaled units of the pentagonal-pyramid and the octahedron configurations are, respectively,
$V^{pyramid}_{6}(\rho_{1},z_{1})=$ $\;\frac{5}{2}\rho_{1}^{2}+15\alpha^{2}z^{2}_{1}+\frac{\sqrt{25+10\sqrt{5}}}{\rho_{1}}+\frac{5}{\sqrt{\rho^{2}_{1}+36z^{2}_{1}}},$ (S6)
$V^{octahedron}_{6}(\rho_{2},z_{2})=$  $\;2\rho_{2}^{2}+\alpha^{2}z^{2}_{2}+\frac{1+2\sqrt{2}}{\rho_{2}}+\frac{8}{\sqrt{\rho_{2}^{2}+z^{2}_{2}}}+\frac{1}{2z_{2}}.$ (S7)
where the coordinates ($\rho_{1}$, $z_{1}$) and ($\rho_{2}$, $z_{2}$) are depicted in Fig. SF1. The minima of these functions intersect at $\alpha\approx 1.238$.
II Mode softening of the 5-ion pentagon
II.1 Out-of-plane normal modes
@@
Figure SF2: The out-of-plane collective oscillations of the 5-ion planar pentagon.
@@
where $l$ is an integer characterizing the angular dependence. These modes are sketched in Fig. SF2: c.o.m. Z-Motion for Cosine l=0 mode, rigid tilt/rotation about Y-axis for Cosine l=1 mode, rigid tilt/rotation about X-axis for Sine l=1 mode, and two buckling modes for Cosine and Sine l=2.
II.2 Buckling critical point